OpenStudy (anonymous):

Find the range of y equals 2 divided by x, plus 2.

OpenStudy (anonymous):

a fraction is only zero if the numerator is zero your range will be all real numbers except 0, because your numerator is 2

OpenStudy (anonymous):

A. All Real Numbers except y = 2 B. y ≥ 2 C. y ≤ 2 D. y ≥ 0 So which would it be? I choose D and got it wrong.

OpenStudy (skullpatrol):

$$\huge y=\frac{2}{x}+2$$

OpenStudy (anonymous):

Yeah that is my equation.

OpenStudy (anonymous):

oh sorry i thought it was \[y=\frac{2}{x+2}\]

OpenStudy (skullpatrol):

Start by plotting values :-)

OpenStudy (skullpatrol):

You know x cannot equal 0, right?

OpenStudy (anonymous):

and more to the point \(\frac{2}{x}\) cannot be zero

OpenStudy (anonymous):

so if \[\frac{2}{x}\] cannot be zero, what can't \(\frac{2}{x}+2\) be?

OpenStudy (anonymous):

All Real Numbers except y = 2

OpenStudy (skullpatrol):

Yes.